OpenStudy (anonymous):

if dy/dx= (x^2)(y^2 ), then d^2y/dx^2 = .... Not sure if I should use differiantial equation to solve or product rule. Help please!!!

zepdrix (zepdrix):

\[\large y'=x^2 y^2\] Find \(\large y''\) Product rule \c:/

OpenStudy (anonymous):

Do you know the final answer?

zepdrix (zepdrix):

Do I know the answer? :O yes.. but try to work it out for yourself silly! Here is the setup for product rule. \[\large y''=\color{royalblue}{(x^2)'}y^2+x^2\color{royalblue}{(y^2)'}\] Need to take the derivative of the blue terms.

OpenStudy (anonymous):

I got 2xy^2/ 1-(2y)(x^2) = dy/dx

zepdrix (zepdrix):

I'm not sure why you have a divide by 1 in the middle, that's kinda confusing. Is this what you came up with? We'll need to make a minor fix if so. \[\large y''=2xy^2-2yx^2\]

OpenStudy (anonymous):

yes that is what I got when I did product rule.

OpenStudy (anonymous):

But don't we have write dy/dx for the derivative of 2y?

zepdrix (zepdrix):

Yes very good, we have to multiply the 2y by dy/dx. Since we took our derivative `with respect to x`, Whenever we take the derivative of y, we have to connect a dy/dx term to it.

zepdrix (zepdrix):

\[\large y''=2xy^2+2yx^2y'\] Another small error, it should be addition in the middle. I'm not sure why it changed to subtraction when you took the derivative. hmm

zepdrix (zepdrix):

Depending on what your teacher wants ~ You may be asked to simplify this down a little bit further. Remember that our y' was given to us to be,\[\large \color{orangered}{y'=x^2 y^2}\] We can substitute this in for our y' that showed up if we want to simplify it down further. \[\large y''=2xy^2+2yx^2\color{orangered}{y'}\]

OpenStudy (anonymous):

got it! 2x^4y^3+ 2xy^2

zepdrix (zepdrix):

yay good job! c:

OpenStudy (anonymous):

Thanks.